Tom bought a computer for 15% off from the list price of p dollars. If the sales tax was 8%, how much did he pay for the computer including sales tax?
Answer by Theo(13342)


list price = p

sale price = p - .15 * p = (1 - .15) * p = .85 * p.

tax on sale price = .08 * .85 * p = .068 * p.

total price = sale price + tax on sale price becomes:

total price = .85 * p + .068 * p = (.85 + .068) * p = .918 * p.

he paid .918 * p for the computer.

this price included the tax on the sale price.

for example:

if the list price was 250 dollars, then he paid .918 * 250 = 229.5.

the list price was 250.
the sale price was .85 * 250 = 212.5
the tax on the sale price was .08 * 212.5 = 17
the total price was the sale price plus the tax on the sale price = 212.5 + 17 = 229.5.